OpenStudy (gm100299):

little help please? consider this function in recursive form f(1)=-2 f(n)=4f(n-1) where n>/- 2 select the equivalent explicit function for n >/- 1 a. f(n)=-2n b. f(n)=-2(n-1) c. f(n)=-2(4)^n-1 d. f(n)=-2(n)^4

OpenStudy (holsteremission):

You can find the general rule for \(f(n)\) by considering one value of the function at a time. Start by finding \(f(2)\). According to the formula, you should have \[f(2)=4f(2-1)=4f(1)\] Next, find \(f(3)\). Well, you have again \[f(3)=4f(3-1)=4f(2)=4(4(f(1))=4^2f(1)\] Continue on and you should notice a distinct pattern. For any value of \(n\), you can write \(f(n)\) as a function of \(f(1)\) times some power of \(4\). What is the pattern?